### v1

@@ 2.2 Human Benchmarks
Any Property Damage or Injury Scanlon et al. (2023) Any Property Damage or Injury Blanco-Adjusted
21.0
22.3
17.9
15.3
2,109,149
24,224
927
28,445
Mileage Blended N/A 21.2
18,530
Flannagan et al. (2023) N/A N/A
64.9
N/A
@@
Police-Reported Scanlon et al. (2023) Police-Reported Unadjusted
4.31
4.55
5.77
3.71
2,109,149
24,224
927
28,445
Mileage Blended N/A 4.84
18,530
Any-Injury-Reported Scanlon et al. (2023) Any-Injury-Reported Unadjusted
1.22
1.29
3.79
1.65
2,109,149
24,224
927
28,445
Mileage Blended N/A 1.91
18,530
Scanlon et al. (2023) Any-Injury-Reported Blincoe-Adjusted
1.78
1.88
5.55
2.41
2,109,149
24,224
927
28,445
Mileage Blended N/A 2.78
18,530
2.3 Confidence Intervals
@@
Two types of confidence intervals are presented in this paper. First, confidence intervals are estimated for Incidents per Million Miles (IPMM) using a Poisson exact model. The confidence interval, IPM, can be computed as
@@
where $qgamma$ is the quantile function of the gamma distribution (i.e., inverse cumulative distribution function), $\alpha$ is the significance level (e.g., 0.05), $n$ is the event count, and $m$ number of miles driven. Second, confidence intervals for the ratio between the ADS crash rate of $Y$ events over $t$ miles and the benchmark crash rate of $X$ events over $s$ miles were calculated using the method described in Nelson (1970) as
@@
where $qbetaprime$ is the quantile function of the beta prime distribution (i.e., inverse cumulative distribution function). This formulation of the confidence intervals can be computed in the event of zero event counts in the ADS data (i.e., $Y=0$), but as a result may be more conservative (i.e., wider confidence intervals) in the case of non-zero ADS events counts compared to the Poisson exact model.
Another interpretation of the ratio between ADS rate ($\mu=Y/t$) and benchmark rate ($\lambda=X/s$) is as a relative difference in the rate, $\delta$,
@@
Because this percent change in the rate is a direct product of the ratio, the same confidence intervals described in Equation 2 can be used to generate confidence intervals in the relative difference.
3 Results
In total, this study considered the first 7.14 million miles of driving by the Waymo RO service, which drove 5.34 million miles in Phoenix, 1.76 million miles in San Francisco, and 0.0467 million (46k) miles in Los Angeles. Table 3 shows the number, IPMM, and 95% confidence intervals for the different event measures examined in this study. In Los Angeles, there was a single event that was an SGO-Reported crash that also met the in-transport and exclude low delta-V inclusion criteria resulting in an IPMM of 21.4 with 95th percentile confidence intervals of (0.5, 119) IPMM. There were no (zero) police-reported and any-injury-reported crashes in Los Angeles, resulting in a 0 IPMM with 95th percentile confidence intervals of (0, 64.1) IPMM.
@@ 3 Results
RO Miles (millions) 5.33 - - 1.75 - -
SGO-Reported 38 7.1 (5.0, 9.8) 33 18.9 (12.9, 26.4)
SGO-Reported in Transport 33 6.2 (4.3, 8.7) 29 16.6 (11.1, 23.7)
SGO-Reported Exclude Low Delta-V 18 3.4 (2.0, 5.3) 14 8.0 (4.4, 13.4)
SGO Police-Reported 12 2.2 (1.2,3.9) 3 1.7 (0.4,0.5)
@@
Table 4 reports the rate ratio of ADS to Human crashed vehicle rates and their 95th percentile confidence intervals for property damage crashes. Results with bold print and an asterisk indicate statistical significance (i.e., the 95th percentile confidence intervals do not cross a 1.0 rate ratio). Compared to the benchmarks derived from human crash reports adjusted for underreporting, the ADS crashed vehicle rate is lower in Phoenix, San Francisco, and when considering all locations together, while higher in Los Angeles. Only the Phoenix and aggregate comparisons are statistically significant to a 95% level for both the all NHTSA SGO-Reported and SGO-Reported Excluding Low Delta-V ADS outcome groups.
The result was statistically significant in San Francisco for the SGO-Reported Excluding Low Delta-V group only. Compared to the NDS benchmarks, the ADS also has a lower rate compared to humans, which was statistically significant. The magnitude of the difference between the ADS and the human benchmarks varies depending on the estimate type (underreporting adjusted or observed) and with the exclusion of low delta-V events in the ADS data. As stated above, there was 1 crash in Los Angeles which met both the SGO and low delta-V thresholds. Compared to the Blanco-adjusted benchmark, the rate ratio was 1.40 with confidence intervals of (0.02, 8.92), which indicates the result is not statistically significant. As discussed in the methods section, the large discrepancy between the NDS studies and adjusted any property damage or injury crashed vehicle rates suggest that there is systematic uncertainty in the data, which makes interpreting the effect of the ADS on any property damage or injury outcomes difficult.
@@
Table 4: Comparison of Waymo ADS and Human Benchmark Crashed Vehicle Rate for Crashes with Any Property Damage or Injury.
All NHTSA SGO-Reported SGO-Reported Excluding Low Delta-V (<1 mph)
@@
Human Benchmark Location Human IPMM ADS IPMM ADS to Human Rate Ratio 95th Percentile Confidence Intervals ADS IPMM ADS to Human Rate Ratio 95th Percentile Confidence Intervals
Any Property Damage or Injury Crashes-Blanco Adjustment (Scanlon et al., 2023)
Phoenix
22.3
6.2
0.28*
0.18 0.41
3.4
0.15* 0.08 0.25
San Francisco
17.9
16.5
0.92
0.58
1.39
8.0
0.45* 0.22 0.80
Total - Mile Blend
21.2
8.8
0.42* 0.31 0.55
4.5
0.21* 0.14 0.31
Total - National Average
21.0
8.8
0.42* 0.31 0.55
4.5
0.21* 0.14 0.32
UMTRI/VTTI Ride-Hailing (Flannagan et al., 2023)
San Francisco
64.9
18.8
0.29* 0.19 0.43
8.0
0.12* 0.06 0.22
SHRP-2 NDS (Blanco et al., 2016)
All Locations
20.1
10.1
0.50* 0.37 0.66
4.5
0.22* 0.14 0.33
@@
Table 5 shows the ADS to Human rate ratio and their 95th percentile confidence intervals for police-reported and any-injury-reported crashed vehicle rates. The ADS had a lower police-reported and any-injury-reported crashed vehicle rate that was statistically significant in Phoenix, San Francisco, and when considered in aggregate. There were no observed police-reported or any-injury-reported crashes in Los Angeles resulting in an rate ratio of 0. The result in Los Angeles was not statistically significant due to low mileage (confidence intervals of [0, 21.3] for police-reported and [0, 32.8] for any-injury-reported). Inverting the rate ratios in Table 5, humans have between 2.0 and 3.3 times the rate of police-reported crashes and between 4.2 and 10 times the rate of any-injury-reported crashes than the Waymo ADS. When considering all locations together compared to the mileage blended benchmark, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.78 IPMM for the human benchmark, an 85% reduction or a 6.8 times lower rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.85 IPMM for the human benchmark, a 57% reduction or 2.3 times lower rate.
@@
ADS Events Human Benchmark Location Benchmark IPMM ADS IPMM ADS to Human Rate Ratio 95th Percentile Confidence Intervals
Police-Reported Observed police reported (Scanlon et al., 2023)
Phoenix
4.55
2.2
0.49*
0.23 0.92
San Francisco
5.77
1.7
0.30* 0.05 0.96
Total Mileage Blended
4.85
2.1
0.43* 0.22 0.76
Total National Average
4.31
2.1
0.49* 0.25 0.85
Any Injury- Reported Any Injury (with underreport adjustment) (Scanlon et al., 2023)
Phoenix
1.88
0.4
0.19* 0.02 0.81
San Francisco
5.55
0.6
0.10* 0.001 0.66
Total Mileage Blended
2.78
0.4
0.15* 0.02 0.49
Total National Average
1.78
0.4
0.24* 0.04 0.77
@@
Figure 2 shows the percent reduction of the ADS for police-reported and any-injury-reported crashed vehicle rates and their 95th percentile confidence intervals relative to the human benchmarks from Scanlon et al. (2023). The ADS reduced police-reported crashes between 51% and 70% and any-injury-reported crashes (adjusted for underreporting) between 76% and 90%. A comparison to the observed any-injury-reported rates from Scanlon et al. (2023) are shown in Figure 2 as a demonstration of the sensitivity of the underreporting adjustment to the results. As mentioned in the methods section, an underreporting adjusted any-injury-reported rate is the most valid comparison to ADS crashes due to known underreporting of injuries in human crash data, while there is low likelihood of injury underreporting in ADS data. In general, the same conclusions are drawn by comparing the ADS observed and adjusted any-injury-reported rates, except for in Phoenix and when compared to a national benchmark where the confidence intervals overlap 0%. Although the point estimates are still lower, because the benchmark rate and observed ADS crash count is small, a small decrease in the benchmark rate greatly increases the confidence intervals.

### v2

@@ METHODS
Any Property Damage or Injury Scanlon et al. (2023) Blincoe-adjusted Any Property Damage or Injury 9.40 9.43 10.49 6.84
2,140,140 24,865 862 29,952
Flannagan et al. (2023) Ride-hailing NDS N/A N/A 64.9 N/A
@@
Police-Reported Scanlon et al. (2023) Observed Police-Reported 4.10 4.31 5.86 3.39
2,140,140 24,865 862 29,952
Any-Injury-Reported Scanlon et al. (2023) Observed Any-Injury-Reported 1.21 1.24 3.98 1.54
2,140,140 24,865 862 29,952
Scanlon et al. (2023) Blincoe-adjusted Any-Injury-Reported 1.76 1.81 5.82 2.26
2,140,140 24,865 862 29,952
This study used two types of data sources to match Any Property Damage or Injury benchmarks to the ADS data. First, police-reported crash data from national (the NHTSA CRSS and Fatality Analysis Reporting System, FARS) and state crash databases were adjusted for human underreporting in Scanlon et al. (2023) using underreporting adjustments for property damage only and injury crashes from Blincoe et al. (2023). The Blincoe et al. (2023) underreporting adjustments were developed using a combination of national police-report data, insurance claims data, and a nationally representative phone survey. These national underreporting estimates for property damage only and injury crashes were then applied to national and state crash data from 2022 in Scanlon et al. (2023). The state data was restricted to the counties where ADS deployments are currently occurring (San Francisco, CA, Maricopa, Arizona, and Los Angeles, CA). Both the state and national data were also restricted to crashes involving passenger vehicles and occurring on surface streets to match the operating conditions of the ADS. The human VMT data in Scanlon et al. (2023) was sourced from Federal Highway Administration (FHWA) for national and state data using the same conditions (passenger vehicles on surface street roads). Second, observed crash rates from NDS were used. Flannagan et al. (2023) derived crash rates from a fleet of vehicles mostly used for ride-hailing on surface streets in San Francisco; these rates were then compared to ADS crashes in San Francisco. Blanco et al. (2016) derived an crash rate using the SHRP-2 NDS, which were personally-owned passenger vehicles that were operated on all road types.
For the comparison of Waymo’s RO crash rate to available human benchmark crash rates, the following subsets of Waymo’s SGO-reported crashes were identified: (1) Those crashes reported by Waymo and without an autonomous specialist behind the wheel (RO crashes), (2) in-transport vehicles (defined as crashes where the Waymo vehicles was not parked in a parking spot), and (3) those where the Waymo vehicle was not impacted (defined as those crashes where the Waymo vehicle struck or was struck by another road user or object). In those ADS crashes with an autonomous specialist seated in the driver’s seat, it is difficult to disambiguate the contributions of the ADS and the autonomous specialist. Therefore, this analysis excluded those crashes with an autonomous specialist present behind the wheel, as RO crashes are most representative of the intended autonomous ride-hailing service. The crashed vehicle count from the police reported benchmark data included only in-transport, impacted vehicles. Therefore, the any property damage or injury benchmarks from Scanlon et al. (2023) were compared with all SGO-reported ADS crashes, except those where the ADS vehicle was not-in-transport or the ADS vehicle was not impacted. Because the NDS data sources described in Flannagan et al. (2023) and Blanco et al. (2016) did not specify whether not in-transport NDS vehicles were excluded, we compared these NDS data to all reported SGO ADS crashes excluding when the ADS vehicle was not impacted. All SGO collisions that met the conditions were included in the study, regardless of crash mode or contacted object.
In addition to examining all NHTSA SGO-reported crashes, results were also reported for a dataset that excluded minor crashes, which we will refer to as “SGO Crash Excluding Low Delta-V.” Although the benchmark data was either adjusted for underreporting or used NDS data designed to find most contacts, there is still uncertainty about the lower reporting threshold in the benchmark data and underreporting corrections. This result of excluding low delta-V ADS crashes was presented alongside the larger all SGO-reported crashes group to investigate the sensitivity of lower reporting threshold on the comparison to the benchmark. We defined these minor contacts as those in which a crash reconstruction model found that both vehicles had a change in velocity (i.e., a delta-V) of less than 1.0 mph. Crash delta-V was determined by a crash model (for vehicle-to-vehicle crashes) or estimated by collision reconstruction techniques (for fixed object crashes).
The human crash benchmarks reported in Scanlon et al. (2023) for Phoenix, San Francisco, and Los Angeles were combined proportional to the miles driven in the Waymo RO service, as shown in Table 3. This mileage blended average was compared with the ADS crashed vehicle rate aggregated in the three locations. For the purposes of confidence intervals, the mileage blended benchmark used 19,002 million VMT, which is the average benchmark VMT proportional to the miles driven in the Waymo RO service.
Table 3: Waymo Mileage Blended Crashed Vehicle Rates and VMT for Human Benchmarks from Scanlon et al. (2023)
Outcome Group Human Benchmark Source Human Benchmark IPMM
Any Property Damage or Injury Blincoe-Adjusted 9.67
Police-Reported Observed Police-Reported 4.68
Any-Injury-Reported Observed 1.92
Blincoe-Adjusted 2.80
The second comparison made in this study was between ADS NHTSA SGO crashes with a police report filed and the observed human police-reported rate. Police-reported and alleged-injury crashes were identified from the SGO-reported crashes using fields in the NHTSA SGO data. The human police-reported rates in Scanlon et al. (2023) did not have an underreporting correction applied. Third, ADS NHTSA SGO crashes where any injuries were reported were compared to the human injury benchmark from Scanlon et al. (2017), both adjusted for underreporting using Blincoe et al. (2023) and unadjusted. Because police-reported and any-injury-reported crashes are relatively rare, most human NDS-based benchmarks have few of these crashes and do not attempt to estimate rates. Therefore, the only comparable benchmark data sources available were the police reported and any injury reported rate estimates presented in Scanlon et al. (2023).
This study did not compare the Waymo ADS performance to all human benchmarks developed in Scanlon et al. (2023). Scanlon et al. (2023) also presented an any property damage or injury crash rate using a property damage only underreporting estimated from the SHRP-2 NDS developed in Blanco et al. (2016). This current study did not compare the ADS to this Blanco-adjusted benchmark because Blanco et al. (2016) only attempted to estimate unreported crashes (those that met the reporting threshold but were not reported) where the Blincoe et al. (2023) underreporting amount accounted for both unreported crashes and crashes that did not meet the reporting threshold (Scanlon et al., 2023). Scanlon et al. (2023) also presented a tow-away benchmark (i.e., any vehicle was towed from the scene due to damage). This study did not compare the ADS to this tow-away benchmark because ADS vehicles are likely towed from the scene more often than human vehicles for reasons that are unrelated to the amount of damage caused by the crash (e.g., due to damage to sensor covers). Finally, the ADS was not compared to the airbag deployed, suspected serious injury+, or fatal injury benchmarks because there are not yet sufficient miles in the ADS data to draw conclusions about these more rare outcomes (see statistical power analysis of Kalra and Paddock (2016) and Scanlon et al. (2023), discussed further in Appendix A.4 of the current paper)
Details of the ADS data selection criteria, the collision model, and a list of all crash incidents included for analysis in this paper is included in Appendix A.1.
Confidence Intervals
@@
Two types of confidence intervals are presented in this paper. First, confidence intervals are estimated for incidents per million miles (IPMM) using a Poisson exact model. The confidence interval, IPMM, can be computed as
@@
where $qgamma$ is the quantile function of the gamma distribution (i.e., inverse cumulative distribution function), $\alpha$ is the significance level (e.g., 0.05), $n$ is the event count, and $m$ is the number of miles driven. Second, confidence intervals for the ratio between the ADS crash rate of $Y$ events over $t$ miles and the benchmark crash rate of $X$ events over $s$ miles were calculated using the method described by Nelson (1970) as
@@
where $qbetaprime$ is the quantile function of the beta prime distribution. This formulation of the confidence intervals can be computed in the event of zero event counts in the ADS data (i.e., $Y=0$), but as a result may be more conservative (i.e., wider confidence intervals) in the case of non-zero ADS events counts compared to the Poisson exact model. Appendix A.3 provides example code and test cases.
Another interpretation of the ratio between ADS rate ($\mu=Y/t$) and benchmark rate ($\lambda=X/s$) is a relative difference in the rate, $\delta$,
@@
Because this percent change in the rate is a direct product of the ratio, the same confidence intervals described in Equation 2 can be used to generate confidence intervals for the relative difference.
The national crash rates from Scanlon et al. (2023) were derived from crash counts from the CRSS database, which features a complex survey design for sampling crashes. Standard errors for counts that use the survey design variables are larger than standard errors calculated using conventional methods using just the weighted counts (Zhang and Diaz, 2020). To investigate the effect of the CRSS variance estimates on the confidence intervals for the rate ratio, the authors constructed confidence intervals using a parametric bootstrap using the standard error for the benchmark crash counts estimated using the survey design variables. The rate ratio confidence intervals computed using the parametric bootstrap method were narrower than those computed using the method described in Equation 2. This result suggests that the method from Nelson (1970) is more sensitive to the other parameters in the rate ratio calculation (i.e., the count of ADS events) than a parametric bootstrap method. The Nelson (1970) method is more conservative (i.e., produces larger confidence intervals) relative to the parametric boostrap method. For this reason, the confidence intervals described in Equation 2 were used to construct rate ratio confidence intervals for the national benchmark comparisons.
RESULTS
In total, this study considered the first 7.14 million miles of driving by the Waymo RO service, which drove 5.34 million miles in Phoenix, 1.76 million miles in San Francisco, and 0.0467 million (i.e., 46.7 thousand) miles in Los Angeles. Table 4 shows the number, IPMM, and 95% confidence intervals for the different event measures examined in this study. Results for Los Angeles, which are based on a single collision in the any property damage or injury category, are listed in Appendix A.2.
@@ RESULTS
RO Miles (millions) 5.34 - - 1.76 - -
SGO-Reported 38 7.1 (5.0, 9.8) 34 19.4 (13.4, 27.1)
SGO-Reported in Transport 33 6.2 (4.3, 8.7) 29 16.5 (11.1, 23.7)
SGO-Reported Exclude Low Delta-V 17 3.2 (1.9, 5.1) 14 8.0 (4.4, 13.4)
SGO Police-Reported 12 2.2 (1.2,3.9) 3 1.7 (0.4, 5.0)
@@
Table 5 and 6 report the rate ratio of ADS to human crashed vehicle rates and 95% confidence intervals for the any property damage or injury outcome group. Results with bold print and an asterisk indicate statistical significance (i.e., the 95% confidence intervals do not contain a 1.0 rate ratio). The human benchmarks are compared to two subsets of the ADS crash data: all NHTSA SGO-reported crashes (Table 5) and SGO-reported crashes excluding low delta-V (< 1 mph) (Table 6). The former is the most broad definition of the any property damage without any lower threshold for property damage. The latter is shown to demonstrate the sensitivity of the results to the minimum inclusion criteria on the ADS side, as the human benchmarks all had different inclusion criteria for any property damage or injury. Of the 63 SGO-reported in transport crashes, 32 crashes (or 51%) had a delta-V greater than 1 mph.
@@
Table 5: Comparison of Waymo ADS and Human Benchmark Crashed Vehicle Rate for Crashes with Any Property Damage or Injury - All NHTSA SGO-Reported Crashes.
@@
Human Benchmark Location Human IPMM ADS IPMM ADS to Human Rate Ratio 95% Percent Confidence Intervals
Scanlon et al. (2023) Blincoe-adjusted Phoenix 9.43 6.2 0.65* 0.43 0.96
San Francisco 10.5 16.5 1.57 0.99 2.37
Total - Mile Blend 9.67 8.8 0.91 0.67 1.20
Total - National Average 8.94 8.8 0.99 0.72 1.30
Flannagan et al. (2023) Ride-hailing NDS San Francisco 64.9 19.4 0.30* 0.19 0.44
Blanco et al. (2016) SHRP-2 NDS All Locations 20.2 10.2 0.51* 0.38 0.67
Table 6: Comparison of Waymo ADS and Human Benchmark Crashed Vehicle Rate for Crashes with Any Property Damage or Injury - SGO-Reported Crashes Excluding Low Delta-V.
Human Benchmark Location Human IPMM ADS IPMM ADS to Human Rate Ratio 95% Confidence Intervals
Scanlon et al. (2023) Blincoe-adjusted Phoenix 9.43 3.2 0.34* 0.18 0.57
San Francisco 10.5 8.0 0.76 0.38 1.36
Total - Mile Blend 9.67 4.5 0.46* 0.30 0.68
Total - National Average 8.94 4.5 0.50* 0.32 0.74
Flannagan et al. (2023) Ride-hailing NDS San Francisco 64.9 8.0 0.12* 0.06 0.22
Blanco et al. (2016) SHRP-2 NDS All Locations 20.2 4.5 0.22* 0.14 0.33
@@
Table 7 shows the ADS to Human rate ratio and 95% confidence intervals for police-reported and any-injury-reported crashed vehicle rates. The ADS had a lower police-reported and Blincoe-adjusted any-injury-reported crashed vehicle rate that was statistically significant in Phoenix, San Francisco, and when considered in aggregate. The observed any-injury-reported comparison was not statistically significant in Phoenix and when compared to a national average, but significant in San Francisco and in the mileage blended comparison. The comparison of the ADS to the observed any-injury-reported benchmark serves as an analysis of the sensitivity of the results to the Blincoe injury underreporting estimate. The results show that even when not accounting for known underreporting in human data, the point estimates still show the ADS had a lower any-injury-reported crash rate, although not all comparisons were statistically significant due to the low event counts and mileage. Inverting the rate ratios in Table 7, humans have between 1.9 and 3.4 times the rate of police-reported crashes and between 4.2 and 10 times the rate of Blincoe-adjusted any-injury-reported crashes as the Waymo ADS. When considering all locations together, the any-injury-reported crashed vehicle rate was 0.41 incidents per million miles (IPMM) for the ADS vs 2.80 IPMM for the human benchmark, an 85% reduction or a human crash rate that is 6.7 times higher than the ADS rate. Police-reported crashed vehicle rates for all locations together were 2.1 IPMM for the ADS vs. 4.68 IPMM for the human benchmark, a 55% reduction or a human crash rate that was 2.2 times higher than the ADS rate. Figure 3 shows the percent reduction of the ADS for police-reported and any-injury-reported crashed vehicle rates and 95% confidence intervals. This percent reduction is another common form of the rate ratios shown in Table 7.
@@
ADS Events Human Benchmark Location Human IPMM ADS IPMM ADS to Human Rate Ratio 95% Confidence Intervals
Police-Reported Scanlon et al. (2023) Observed Police-Reported Phoenix 4.31 2.2 0.52* 0.24 0.97
San Francisco 5.86 1.7 0.29* 0.05 0.95
Total Mileage Blended 4.68 2.1 0.45* 0.23 0.78
Total National Average 4.1 2.1 0.51* 0.26 0.90
Any-Injury-Reported Scanlon et al. (2023) Observed Any-Injury-Reported Phoenix 1.24 0.4 0.30 0.03 1.23
San Francisco 3.98 0.6 0.14* 0.002 0.91
Total Mileage Blended 1.91 0.4 0.22* 0.03 0.71
Total National Average 1.21 0.4 0.35 0.05 1.13
Scanlon et al. (2023) Blincoe-adjusted Any-Injury-Reported Phoenix 1.81 0.4 0.21* 0.02 0.84
San Francisco 5.82 0.6 0.10* 0.001 0.62
Total Mileage Blended 2.80 0.4 0.15* 0.02 0.49
Total National Average 1.76 0.4 0.24* 0.04 0.77